OpenStudy (anonymous):

Find the distance between point (-3,4) and (6,1) and please explain.

OpenStudy (anonymous):

distance = sq rt (x2 - x1)^2 + (y2 - y1)^2 (-3,4) x1 = -3 and y1 = 4 (6,1) x2 = 6 and y2 = 1 now lets sub d = sq rt (6 - (-3)^2 + (1 - 4)^2 d = sq rt (6 + 3)^2 + (-3)^2 d = sq rt (9^2) + (-3)^2 d = sq rt (81 + 9) d = sq rt 90 d = 9.48